OpenStudy (anonymous):

One number is 7 more than another number. The sum of the numbers is 63. What are the numbers?

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + (x + 7) = 63}\) Solve for x, and x + 7.

OpenStudy (anonymous):

x=y+7 x-y=7 x+y=63 add both nd find the answer

OpenStudy (anonymous):

Let the other number be : x So, First number = x + 7 And given: x + (x + 7) = 63 2x = 56 Solve for x..

OpenStudy (anonymous):

that is 28 and 35 28+35=63

OpenStudy (anonymous):

what about my answer?

OpenStudy (callisto):

You've got your answer already...

OpenStudy (anonymous):

some number = x some number +7 = x+7 now combine them to equal to 63 and solve for x. so we have, x+(x+7) = 63 x = 28 now since you know the value of one x the other would be 28 + 7 = 35